Jake’s Cheese Company produces gourmet cheese for resale at local grocery stores. The master budget indicates that the company expects to use 4.0 pounds of direct materials for each unit produced at a cost of $7.00 per pound (one unit = one batch of cheese). Each unit produced will require 0.50 direct labor hours at a cost of $12.00 per hour. Variable manufacturing overhead is applied based on direct labor hours at a rate of $6.00 per hour. Last year's sales were expected to total 50,000 units. Jake just received last year's actual results showing sales of 45,000 units. Required: (1) Calculate the standard cost per unit for direct materials, direct labor, and variable manufacturing overhead. (2) Prepare a flexible budget based on actual sales for direct materials, direct labor, and variable manufacturing overhead.
Answer #1

Standard Cost per Unit for direct materials= 7/4 = 1.75 Standard Cost per Unit for direct Loabour=0.5*12 = 6 Standard Cost pe